OpenStudy (anonymous):

help ill give medal and fan An equation is shown below: 9x - y = -2

OpenStudy (solomonzelman):

What do you want to do with it?

OpenStudy (anonymous):

Part A: Explain how you will show all of the solutions that satisfy this equation. Part B: Determine three different solutions for this equation. Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent.

OpenStudy (solomonzelman):

For part A: A solution to this equation is any point that is on a line (and there are infinity of them). you can plug an x-value (call it "c") and get some value for y, (call it "d") then your solution is (c,d). The reason there are infinity of such solutions, is because the domain is inifinite - you can plug any x-value (as long as it is a real number).

OpenStudy (solomonzelman):

Part B: There are just going to be 3 different examples of such (c,d) solutions. Whatever you plug in for x (to get y). What you are plugging in for x, is absolutely up to you.

OpenStudy (anonymous):

i appreciate yall help

OpenStudy (solomonzelman):

Part C: Inconsistent system of equations can be any y=mx+b y=mx+c (two parallel lines) (where b doesn't equal c) there is no solution that can satisfy them both at the same time. Because: 1) Solution of a system of equations is the point(s) of intersection (of the 2 functions) however parallel lines do not intersect with each other. (Right?) 2) If you subtract the equations, you will get y=mx+b \(\LARGE\color{black}{ ^{-} }\)y=mx+c ------------ 0 = 0 + b - c and thus you get 0=b-c, and b-c is not equal to zero, because b and c are not same numbers. (if they are, then it is same line, not 2 different paralle lines.)

OpenStudy (solomonzelman):

You can just change the "b" of the y=mx+b in your equation, to any other number. And write your old and new (made up by you) equation together. This will answer Part C, getting you an INCONSISTENT system of equations.

OpenStudy (anonymous):

first you gotta a burn a blunt

OpenStudy (anonymous):

im sorry what